RICARDO IS 4 YEARS LESS THAN TWICE AS OLD AS HIS SISTER.
THE SUM OF THEIR AGES IS 20.
HOW OLD IS RICARDO?

Answer by elima(1433)

RICARDO IS 4 YEARS LESS THAN TWICE AS OLD AS HIS SISTER.
THE SUM OF THEIR AGES IS 20.
r=ricardo's age
x=sisters age
r=2x-4
===================
x+2x-4=20
3x=24
x=8
so his sister is 8
Ricardo = 2(8)-4=12
:)